The perimeter of a triangle is 9 and is represented by three consecutive numbers. What are the sides of the triangle. I cannot figure it out, I tried everything. Thank you!
Answer by Edwin McCravy(20086)

The perimeter of a triangle is the sum of the sides, and it must be 9.

Consecutive numbers are 3 counting numbers in a row like this

5,6,7 or
4,5,6 or
3,4,5 or
2,3,4 or
1,2,3 

Now let's try those out to see if any of those could be the three sides
and add up to the perimeter of 9

5+6+7 = 18, nope, that's too much. It has to be only 9.
4+5+6 = 15, nope, that's still too much. It has to be only 9.
3+4+5 = 12, nope, that's still too much. It has to be only 9
2+3+4 =  9, WELL!  What do you know!  That's it, by cracky!

Answer: 2, 3, and 4
are consecutive numbers and they add up to 9.

Edwin